OpenStudy (anonymous):

Layla has a jar filled with 108 coins. If the jar contains only quarters and nickels and has a value of $15.80, how many nickels are in the jar?

jimthompson5910 (jim_thompson5910):

Hint: n+d = 108 0.05n + 0.10d = 15.80

OpenStudy (anonymous):

but how do you figure out how many of each there should be? its simple but has me confused I think im thinking about it too ,uch

jimthompson5910 (jim_thompson5910):

n+d = 108 d = 108 - n ------------------------------------------------------- 0.05n + 0.10d = 15.80 5n + 10d = 1580 5n + 10(108-n) = 1580 solve for n

OpenStudy (kropot72):

@jim_thompson5910 The question relates to nickels and quarters.

jimthompson5910 (jim_thompson5910):

oh my bad, misread and thought it was nickels and dimes

jimthompson5910 (jim_thompson5910):

n+q = 108 q = 108 - n ------------------------------------------------------- 0.05n + 0.25q = 15.80 5n + 25q = 1580 5n + 25(108-n) = 1580 Solve for n

OpenStudy (anonymous):

How would I solve for N? Ive been doing this aggisntment for the past 2 hours, my brain is dead

jimthompson5910 (jim_thompson5910):

5n + 25(108-n) = 1580 5n + 25(108)+25(-n) = 1580 5n + 2700-25n = 1580 -20n = 1580 - 2700 -20n = -1120 I'll let you finish n = ???

OpenStudy (anonymous):

56?

jimthompson5910 (jim_thompson5910):

yep